OpenStudy (anonymous):

I need help! PLEASE someone|| Solve: x3 − 2x^2 + 10x + 136 = 0

OpenStudy (anonymous):

put in standard form first

OpenStudy (anonymous):

then see if you can simplify

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

x3 − 2x^2 + 10x + 136 = 0 -2x ^2 + (10x + 3x ) +136 = 0 simplified form − 2x^2 + 13x + 136 = 0 if i did not make a mistake now is it easier to do?

OpenStudy (anonymous):

-4, 8?

OpenStudy (anonymous):

x = -4